OpenStudy (anonymous):

16. Solve the equation. Show your work. 1/a 6/18 Is this right? 1/a 6/18 a = 0.3

OpenStudy (anonymous):

no. If 1/a = 6/18, then a = 18/6. You can flip both fractions over since they're the only thing in the equation

OpenStudy (anonymous):

would i need to flip 1/a as well?

OpenStudy (anonymous):

18/6 = 3